solve for unknown
3(2c-5)=45

Answer by oscargut(2103)

3(2c-5)=45
2c-5 = 45/3
2c-5 = 15
2c = 15+5
2c = 20
c = 20/2
c = 10
Answer: c=10
You can ask me more at: mthman@gmail.com
Thanks